paddy is three times as old as Siobhana. Seven years ago, he was five times as old as Sibohana.How old is Paddy now?
Answer by oberobic(2304)

Start by defining your terms.
p = Paddy's age
s = Sibohana's age
Now we set up the givens:
p = 3s OR Paddy is 3 times Siobhana's age
p -7 = 5(s-7) OR 7 years ago, Paddy was 5 times Siobhana's age
Notice that we show BOTH p-7 and s-7 as their ages 7 years ago.
Substituting the value of p from the first equation into the second:
3s -7 = 5(s - 7)
Multiplying through the right side:
3s -7 = 5s - 35
Subtracting 3s from both sides:
-7 = 2s - 35
Adding 35 to both sides:
-7 + 35 = 28 = 2s
Dividing both sides by 2:
14 = s. Siobhana is 14.
We know from the setup that:
p = 3s, so p = 3(14) = 42. Paddy is 42.
Check your work!
Substitute the values into the second equation.
p - 7 = 5(s - 7)
42 - 7 = 5(14 -7)
35 = 5(7), which is true.